Based on the historical data, the population proportion of the residents who have generators in a subdivision is 80%. The local newspaper surveyed 20 residents and 9 claimed that they have generators. What is the correct value of test statistic?

1
Expert's answer
2022-01-23T16:53:38-0500

Test statistic: "z=\frac{\frac{9}{20}-0.8}{\sqrt{\frac{0.8(1-0.8)}{20}}}=-3.91."